OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2+18x+70=0

OpenStudy (michele_laino):

hint: we can write: \[\large {x^2} + 18x + 70 = {x^2} + 18x + 81 - 81 + 70 = ...?\]

OpenStudy (anonymous):

I'm confused...

OpenStudy (welshfella):

note that x^2 + 18x + 81 can be written as a perfect square

OpenStudy (michele_laino):

that's right! @welshfella

OpenStudy (anonymous):

Is it like this format: (?) (x+ _ )^2 = _ (x- _ )^2 = -

OpenStudy (michele_laino):

\[\large {x^2} + 18x + 81 - 81 + 70 = \left( {{x^2} + 18x + 81} \right) - 81 + 70 = ...?\]

OpenStudy (michele_laino):

hint: what is (x+9)^2=...?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

notice that 81 is a perfect square

OpenStudy (anonymous):

Would it be (x+9)^2=9 ?

OpenStudy (anonymous):

NO wait so x would be 9!

OpenStudy (michele_laino):

after that computation, you should get this: \[\large \left( {{x^2} + 18x + 81} \right) - 81 + 70 = {\left( {x + 9} \right)^2} - 11\]

OpenStudy (anonymous):

I'm still confused